Solve algebraically using only one variable. The perimeter, which is the sum of all sides, of a rectangle is 48 square inches. The length of the rectangle is three more than twice its width. Find the length and width of the rectangle
Answer by checkley75(3666)

2w+2(2w+3)=48
2w+4w+6=48
6w=48-6
6w=42
w=42/6
w=7 answer for the width.
length=2w+3, 2*7+3, 14+3, 17 for the length.
proof
2*7+2*17=48
14+34=48
48=48